Two coils, held in fixed positions, have a mutual inductance of 100 μH. What is the peak voltage in one when a sinusoidal current given by I (t) = (10.0 A) sin
(1 000t) is in the other coil?
1
Expert's answer
2019-11-22T10:55:56-0500


From the conditions of the problem we have

"M=100\cdot 10^{-6}=1\cdot 10^{-4}[H]"

"\omega=1000[rad]"

"I_m=10[A]"

"I=0"

We write the second Kirchhoff law for the amplitude values of current and voltage in the second coil

"I\cdot X_L+I_m \cdot(\omega \cdot M)-U_m=0"

or

"0\cdot X_L+I_m \cdot(\omega \cdot M)-U_m=0"

"I_m \cdot(\omega \cdot M)-U_m=0"

where will we write

"U_m=I_m \cdot(\omega \cdot M)=10\cdot 1000\cdot1\cdot10^{-4}=1[V]"